What is the gravitational attraction between a 75 kg boy (165 lbs) and the 50 kg girl (110 lbs) seated 1 m away in the next desk?
Answer #1

M1 = 75kg

M2 = 50kg

R = 1m

From F = GM1.M2/R^2

=) F = 6.67×10^-11× 75×50/1 = 25.012×10^-8 N

=) Force F = 25.012×10^-8 N